Date:
Monday, September 7, 2015
Telescope:
CTIO Blanco 4-m
Submitter:
Ross Cawthon
Submitter Email:
Observer Support:
Manuel Hernandez
Night Assistant:
Claudio Aguilera
Also Present:
David James, Camila Naverrete
Problems:
None.
Program:
Proposal:
2012B-0001
Target of Oportunity:
No
Time Observed:
3.5
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
1.5
Time lost other:
0.0
Description:
Dark Energy Survey
Primary Investigator:
Josh Frieman
Email:
Institution:
Fermilab
Instrument:
CTIO 4m DECam
Organization:
NOAO
Total Program Time:
5
Proposal:
2015B-0601
Target of Oportunity:
No
Time Observed:
5.0
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Description:
DECam Search for the Elusive Stellar Counterpart of the Magellanic Stream
Primary Investigator:
J. Chaname
Email:
Institution:
Pontificia Universidad Catolica de Chile
Instrument:
CTIO 4m DECam
Organization:
Chile
Total Program Time:
5
Total Times | Time Observed | Time Engineering | Time lost technical | Time lost weather | Time lost other | Total Program Time |
---|---|---|---|---|---|---|
2015-09-07 | 8.5 | 0.0 | 0.0 | 1.5 | 0.0 | 10 |
*Observing Team*
_DES Observers:_ Ross Cawthon, David James
_Telops:_ Manuel Hernandez, Claudio Aguilera
_Remote Support:_ ---
*Observing Plan*
# DES until 11:30 CLT, then community time, then DES starts again 4:30 CLT
# evening Standard stars
# run ObsTac queue
# morning standards
*Conditions*
# Scattered clouds and humidity around 75% to start the night.
# With humidity stabilized around 75%, dome opened at 9:34 CLT. Plan to wait 10 minutes for telescope to equilibrate temperature after not being opened at all last night. See how conditions go.
# Standards began at 10:01, ObsTac started at 10:13
# Early part of night, seeing quite bad as mirror 2 or so degrees warmer than dome. Image seeing often around 2 arcmin.
# Rest of night conditions pretty good, seeing around 1 arcmin. Clear skies.
*Accomplishments*
# Evening standards (taken late, see above), E8-A, SDSSJ2300-0000
# Satellite and aTmCam checked.
# Early night (bad seeing due to dome/telescope temperature) many g and r band widefield images
# SN-E1 done early in the night, while the seeing was bad (>1.7 arcmin on images)
# At end of night (seeing 1-1.4 arcmin): SN-E2, SN-X2, SN-X3 observed
# Morning standards SDSSJ0100-0000, SDSSJ0320-0000, SDSSJ0320-0000
# Weekly defocus/refocus script ran
*Minor Problems and Relevant Information*
# See weather delays above.
*Notes for Data Management*
# Early night images had poor seeing, including SN-E1.